Using the Taylor Rule, find the appropriate Federal Funds rate (FFR).

Assume that the Fed has  target inflation rate of 2% and a target GDP growth rate of 3%. What FFR should the Fed target if the current inflation is 1% and the growth rate is 2%?

Enter the number without the percent sign.



1
Expert's answer
2021-05-12T15:30:55-0400

We know that,


r = p + 0.5 * y + 0.5 * (p - 2) + 2


where, r = Federal funds Rate (FFR)

p = Inflation Rate

y = Real GDP Gap


a) r = ?

p = 2%

y = 3%


r = 2 + 0.5 * 3 + 0.5 * (2-2) +2

r = 2 + 1.5 +2


r = 5.5 %


b) r = ?

p = 1%

y = 2%


r = 1 + 0.5 * 2 + 0.5 * (1-2) +2

r = 1+ 1 - 0.5 + 2


r = 3.5 %